Twice a number plus another number is 4. Their sum is 2. Find the numbers
Answer by ankor@dixie-net.com(22740)

Twice a number plus another number is 4.
2x + y = 4
:
Their sum is 2.
x + y = 2
:
Elimination would work easily here
2x + y = 4
x + y = 2
-------------subtracting eliminates y, find x
x + 0 = 2
x = 2
:
Use the 1st equation to find y, replace x with
2(2) + y = 4
4 + y = 4
y = 4 - 4
y = 0
:
Check in the 2nd equation
2 + 0 = 2